8. (5 points) A chemical analysis of a surface water yields the following data: lonEq. Conc. (mg CaCO3/L) Ca2+ 2+ 75 100 250The solutions are written under the question, but I need to know how the answer was found.
Answer #1

(a) The carbonate hardness is majorly contributed by the species HCO3^-. Hence, the carbonate hardness is 250 mg/L

(b) Ca^2+ exhibits both carbonate and non-carbonate hardness.

i.e. The carnonate hardness of Ca^2+ = 225 mg/L

Mg^2+ and SO4^2- contributes to both carbonate and non-carbonate.

Here, the carbonate hardness of Mg^2+ and SO4^2- = 100 + 75 = 175 mg/L

Here, the carbonate hardness of Ca^2+ = (carbonate hardness + non-carbonate hardness) of Mg^2+ and SO4^2-.

Therefore, non-carbonate hardness = 225 - 175 = 50 mg/L